OpenStudy (samara8954):

Kathleen had a bag of candy-coated chocolates. Of the 60 candies in the bag, 20% were red, 30% were orange, 30% were green, and 20% were blue. Of the blue ones, 10% had a damaged candy coating. Approximately how many of the blue ones had a damaged candy coating? 18 1 6 12

OpenStudy (samara8954):

@Gatsby

OpenStudy (samara8954):

@hurica1n

OpenStudy (samara8954):

@HugoHuynh

OpenStudy (anonymous):

1

OpenStudy (anonymous):

First, determine how many of the candies were blue.

OpenStudy (anonymous):

20% of 60 = 0.2 * 60 = 12

OpenStudy (anonymous):

Next, find how many of the blue ones were damaged.

OpenStudy (anonymous):

10% of 12 = 0.1 * 12 = 1.2

OpenStudy (anonymous):

Approximately 1 of the blue ones had a damaged candy coating

OpenStudy (samara8954):

so what is the answer

OpenStudy (samara8954):

@hurica1n

OpenStudy (anonymous):

1

OpenStudy (anonymous):

could 1 of you help me?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

wat do ya need help with

OpenStudy (anonymous):

here it is

OpenStudy (anonymous):

o.....wow im smart but not that smart sorry

OpenStudy (anonymous):

i think the 1st one is a

OpenStudy (anonymous):

plzzzzzzzz i really want to go to bed. lol ill medal u for trying and fan and thx

OpenStudy (anonymous):

yw and ty

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

yes sorry